I am doing permutation problems using the formula nPr = n!/(n-r)!. The problem I am having difficulty with is nP3 = 24. I can get it this far:
n(n-1)(n-2)...2x1 / (n-3)(n-4)...2x1 = 24
n(n-1)(n-2) = 24 Then I'm not sure how to finish. Please help.

Answer by stanbon(57290)

nP3 = 24
n!/(n-3)! = 24
n(n-1)(n-2)=24
n(n^2-3n+2)=24
n^3-3n^2+2n-24=0
I graphed this and found n=4 is a solution.
============
Checking:
4P3 = 4!/(4-3)! = 4!/1! = 4! = 24
OK
============
Cheers,
Stan H.